Pairs of markings a set distance apart are made on highways so that police can detect drivers exceeding the speed limit. Over a fixed​ distance, the speed R varies inversely with the time T. In one particular pair of​ markings, R is 4343 mph when T is 88 seconds. Find the speed of a car that travels the given distance in 77 seconds.
Requals= ? mph ​(Round to the nearest whole​ number.)

Answer by stanbon(75887)

Pairs of markings a set distance apart are made on highways so that police can detect drivers exceeding the speed limit
. Over a fixed​ distance, the speed R varies inversely with the time T.
R = k/T
-----------------------------
In one particular pair of​ markings, R is 4343 mph when T is 88 seconds.
Solve for "k"::
4343 = k/88
k = 382184
Equation:
R = 382184/T
---------------------
Find the speed of a car that travels the given distance in 77 seconds.
Requals= ? mph ​(Round to the nearest whole​ number.)
R = 382184/77 = 4963.43
--------------
Cheers,
Stan H.
------------